# TransParking: A Dual-Decoder Transformer Framework with Soft Localization for End-to-End Automatic Parking

### v1

@@ Abstract
In recent years, fully differentiable end-to-end autonomous driving systems have become a research hotspot in the field of intelligent transportation. Among various research directions, automatic parking is particularly critical as it aims to enable precise vehicle parking in complex environments. In this paper, we present a purely vision-based transformer model for end-to-end automatic parking, trained using expert trajectories. Given camera-captured data as input, the proposed model directly outputs future trajectory coordinates. Experimental results demonstrate that the various errors of our model have decreased by approximately 50% in comparison with the current state-of-the-art end-to-end trajectory prediction algorithm of the same type. Our approach thus provides an effective solution for fully differentiable automatic parking.
@@ I INTRODUCTION
The traditional automatic parking systems are usually composed of three modules: perception, planning and control. In these systems, the vehicle initially employs sensors such as cameras and LiDAR to detect and model its surrounding environment, identifying both the target parking space and any nearby obstacles. Subsequently, a trajectory is planned based on the acquired perception data, and the vehicle’s speed and direction are continuously adjusted in real time—often via a PID controller or other conventional control algorithms to execute the parking maneuver. Although many of these algorithms have been successfully deployed in real-world vehicles, forming a mature industrial ecosystem, they remain highly dependent on accurate environmental information and precise vehicle dynamics models. As a result, any errors introduced during the perception phase can be amplified as they propagate through the planning and control stages, potentially leading to sensor deviations and the failure of the overall system.
Inspired by the Transformer model[1], this paper proposes a pure vision-based end-to-end autonomous driving approach, effectively mitigating the aforementioned problems. We sample trajectory points at certain intervals from expert trajectories and regard them as discrete tokens with positional encodings, similar to word tokens in natural language processing tasks. Subsequently, we utilize encoder and Dual-decoder structure for autoregressive trajectory prediction, decouple the X and Y coordinates of the historical trajectories, and input them separately into the proposed model. The specific contributions of this work are as follows:
@@
Figure 1: Method Overview: After inputting the images captured by cameras and the coordinates of the target slot marked by the client, a Gaussian kernel distribution map is generated. After convolution, the BEV fusion feature is obtained through the attention mechanism. During the decoding process, Dual-Stream Interactive Self-Attention enables the already separated coordinates to mutually pay attention to the coordinate information of each other in the historical trajectory. Before outputting the final result, Soft Localization Coordinate Refinement Head fine-tunes the output result by utilizing the generated attention map.
@@ II RELATED WORK
II-A BEV view in autonomous driving
Bird’s Eye View (BEV) representations offer a clear, top-down perspective, showing vehicles, roads, and obstacles at their true planar positions. Consequently, BEV is widely adopted in autonomous driving applications. However, cameras cannot directly acquire distance information for each pixel. To address this, LSS [2] compensates for the lack of depth information by predicting the depth distribution of each pixel. BEVDepth [3], on the foundation of LSS, incorporates explicit depth supervision and a Depth Refinement Module, enhancing the reliability of depth estimation. BEVDet[4] was improved on the basis of LSS and achieved higher accuracy. BEVDet4D [5], on the basis of Bevdet, introduces a time dimension, achieving the fusion of spatiotemporal features. BEVFormer[6] is a Transformer-based BEV feature generation algorithm. It can simultaneously integrate the spatial and temporal information of multi-view images and reduce the computational load, generating high-quality features. DualBEV[7] introduces a dual feature fusion strategy to effectively predict BEV probabilities. BEVfusion[8] implements a multimodal fusion perception scheme that combines LiDAR and image data, effectively preserving both geometric and semantic information. FastBEV [9] proposed Fast-Ray, thereby significantly enhancing the inference speed.
@@ II-B End-to-end Autonomous Driving
In recent years, the application of neural networks in autonomous driving has significantly advanced [10]. Overall, auto-driving tasks can be broadly divided into urban driving scenarios and parking scenarios. Below, we review several representative works in each category.
Urban driving scenarios: CIL[11] directly maps front-view images, current control inputs, and high-level commands to control signals, but its performance can be suboptimal during sharp turns. To address this issue, TCP[12] introduces a Trajectory-guided Attention mechanism, ultimately yielding multi-step control predictions based on trajectory forecasting. TransFuser [13] realizes the multimodal fusion of image and LiDAR data through the utilization of the self-attention mechanism. Transfuser++[14] investigates the phenomenon of target point shortcuts and refines the widely used baseline model Transfuser. TAT[15] builds upon Transfuser by incorporating target points and demonstrating the significance of the target point attention mechanism. Pix2Planning [16] proposes a Transformer-based approach, in which the target position and historical trajectories are concatenated as the input sequence for the decoder. Yuandong Lyu et al. [17] fuse the bird’s-eye view features based on cameras and those based on LiDAR, and ultimately output the predicted trajectory points and control signals.
Parking scenarios: Chai Runqi et al. [18] designed and verified a two-stage hybrid optimization framework that effectively combines global intelligent search and local gradient optimization. Peizhi Zhang et al. [19] proposed an end-to-end automatic parking system based on reinforcement learning, which directly maps parking space information to the optimal control instructions of the steering wheel angle. Nevertheless, it is hard to guarantee that the reward function is optimal. ParkPredict[20] introduces a nested CNN-LSTM model for predicting trajectory points. Building on this framework, ParkPredict+[21] replaces the LSTM with a transformer and proposes a CNN-based intention prediction network to forecast future trajectory points. Chai Runqi et al. [22] designed a motion planner based on RDNN and a tracking controller based on ALNN. E2E Parking[23] employs a transformer-based architecture that directly maps images and motion states to parking control signals once a user selects a parking spot. ParkingE2E[24] extends the E2E Parking framework. In this approach, a binary map of the target slot is convolved to generate a query vector, which is subsequently fused with the BEV feature map to predict the target point’s coordinates. This method has been validated through real-world testing and has demonstrated excellent performance.
@@ III-A Problem Description and Optimization Goals
[…]
@@ III-B End-to-End Neural Network Model Based on Camera
Query: In the ParkingE2E[24], the authors first generate a binary map with the same spatial dimensions as the BEV feature map and a depth of 1. In this map, the target slot and its surrounding fixed region are marked with a value of 1, while the area outside this range is marked with 0. After applying convolutional neural network, this map is used as a query vector to compute attention with the BEV feature map, serving as the final output of the encoder. This approach offers us significant inspiration and insight. The method of ParkingE2E clearly indicates to the model which areas need attention, However, a drawback is that it does not distinguish subregions within the target area, and the abrupt boundary transition might adversely affect continuity in attention. Therefore, in our work, we aim to give greater weight to the center while retaining some information from the surrounding areas, and this weight gradually weakens as the distance increases. We can select the desired parking slot by using the interactive tool ‘2D Nav Goal’ of Rviz and send the coordinates to the model. Subsequently, a zero tensor with the same width and height as the BEV feature map and a depth of 1 is created. In this channel, a probability map following a Gaussian kernel distribution is generated within the specified range centered on the coordinates of target slot. Subsequently, the tensor undergoes feature extraction via same convolutional neural network, transforming it into a high-dimensional vector. The formula is as follows:
@@
Key and Value: We employ EfficientNet [25] to extract image features $F_{img}\in\mathbb{R}^{C\times H_{img}\times W_{img}}$ from the RGB input, and then use the LSS[2] method to generate the BEV feature map. First, we predict the depth distribution for each pixel and map it to the 3D space. Then, we fuse the depth information with the image features. Finally, we project the fused features onto the BEV voxel grid using the camera extrinsics and intrinsics to generate the camera features. For the obtained feature map, the range in the $x$-direction is denoted as $[-R_{x},R_{x}]$ m, where m indicates meters, and the range in the $y$-direction is denoted as $[-R_{y},R_{y}]$ m.
@@
We separate the X and Y coordinates of the model in the ParkingE2E [24], and then train and predict using two decoders with the same number of layers. At each time step, the model simultaneously outputs the X and Y coordinates, and they can only see the historical information of the same type of coordinates. Based on the findings of the comparative experiments depicted in Table I, the experimental outcome of jointly determining the most appropriate X-coordinate and Y-coordinate at one time based on the perceived surrounding environmental information surpasses the approach of splitting the X-coordinate and Y-coordinate into two time steps.
@@
To address the deficiency that the X and Y coordinates have no knowledge of each other’s historical information, we propose to initially concatenate the feature vectors of the X and Y coordinates at each time step, then conduct a self-attention mechanism calculation and obtain the fused feature vector. Subsequently, the fused feature vector is separated into x and y, which are respectively input into their respective decoders. The specific structure is presented as shown in the figure 3. This attention module ensures that the high-dimensional feature vectors of the X and Y coordinate at the same time step are capable of integrating each other’s feature information. It guarantees that, in the subsequent mask self-attention layers in their respective decoders, The $X_{t}$ and $Y_{t}$ can indirectly take into account the X-coordinate and Y-coordinate information at the 0 to $t_{-1}$ moment.
@@ III-C Lateral control and longitudinal control
In our work, we adopt the similar approach as ParkingE2E[24] for Lateral and Longitudinal Control of non-omnidirectional vehicle. During the control process, we mark the initial moment of parking as $t_{0}$. The predicted path $\mathcal{X}_{t_{0}},\mathcal{Y}_{t_{0}}=\mathcal{N}_{\theta}(I_{t_{0}},S)$ obtained at the initial moment is used. The relative pose from the initial moment $t_{0}$ to the current moment t can be obtained through the positioning system, denoted as $ego_{t_{0}\rightarrow t}$. The target steering angle $\mathcal{A}^{tar}$ is obtained using the RWF (Rear-wheel Feedback) method, and its expression is as follows:
@@
Figure 5: The inference results in indoor and outdoor scenarios. The orange line represents the inference trajectory of the method proposed in this paper, and the purple line represents that of ParkingE2E [24].
@@ IV-A The construction of the dataset
The dataset of this work originates from the publicly available real-vehicle dataset offered by ParkingE2E[24], which does not contain sudden obstacles and negative samples. This dataset acquires RGB images around the vehicle through on-board cameras and employs sensor data fusion algorithms to achieve precise vehicle positioning. Meanwhile, it provides the two-dimensional coordinates of the parking target slot in the ego coordinate system for each time step.
@@ IV-B Evaluation Metrics
To guarantee the fairness of the comparative experiments, the basic training parameters employed by all the models remain consistent. The size of the BEV Features is $200\times 200$, The actual spatial range is $x\in[-10\mathrm{m},10\mathrm{m}]$, $y\in[-10\mathrm{m},10\mathrm{m}]$ with a resolution of $0.1$ meters. In decoders, the maximum value of the trajectory serialization $N_{t}$ is $1200$. The trajectory decoder generates a sequence of predictions with a length of $30$. Our training equipment: One NVIDIA GeForce RTX 4090 GPU. The experimental outcomes of the ParkingE2E model[24] proposed in previous works, the transfuser model (based on the GRU decoder)[13], and our proposed model are presented in Table II. The data indicates that, under the same dataset and parameters, the model we proposed has the highest accuracy rate, and all errors are reduced by approximately 50% compared to ParkingE2E.
@@
ParkingE2E[24] 0.386 0.1193 2.431
Transfuser[13] 0.5293 0.3517 8.674
Figure 6: The performance of the two algorithms in the same scenario where an obstacle suddenly appears. The orange line represents the inference trajectory of the method proposed in this paper, and the purple line represents that of ParkingE2E[24]. Each row of images shows the trajectory inference results of four time steps from left to right.
@@ IV-D Visualization and Analysis of inference
We visualize the trajectory of model inference in Rviz, as shown in Figure 5. We also demonstrate that in the absence of special scenarios such as the sudden appearance of fast-moving objects in the training set, if other vehicles or pedestrians suddenly approach rapidly during the parking process, the model will immediately change its trajectory to prevent a collision. However, the model of ParkingE2E [24] mainly focuses its attention on the target area, thus being unable to modify the vehicle’s driving trajectory in a timely manner, as shown in Figure 6. Meanwhile, we noticed that during the parking process, if the vehicle is still at a distance from the parking space and there are no obstacles, our model will first plan a relatively straight trajectory to avoid turning too early and crossing the parking line. Only when the vehicle approaches the parking space will it start to plan a curve to achieve the turn. Moreover, with the same number of Predicted token, the trajectory inferred by ParkingE2E tends to have a larger range and sometimes deviates from the route, inferring a tortuous path. The specific effect is shown in Figure 7.
Figure 7: The performance of the two algorithms in the same indoor scene. The orange line represents the inference trajectory of the method proposed in this paper, and the purple line represents that of ParkingE2E[24].

### v2

@@ Abstract
In recent years, fully differentiable end-to-end autonomous driving systems have become a research hotspot in the field of intelligent transportation. Among various research directions, automatic parking is particularly critical as it aims to enable precise vehicle parking in complex environments. In this paper, we present a purely vision-based transformer model for end-to-end automatic parking, trained using expert trajectories. Given camera-captured data as input, the proposed model directly outputs future trajectory coordinates. Experimental results on the ParkingE2E real-vehicle dataset show that our method reduces the Hausdorff distance from 0.386 m to 0.2156 m, the L2 distance from 0.1193 m to 0.06296 m, and the Fourier descriptor difference from 2.431 to 1.239, under identical settings. Our approach thus provides an effective solution for fully differentiable automatic parking.
@@ I INTRODUCTION
The traditional automatic parking systems are usually composed of three modules: perception, planning and control. In these systems, the vehicle initially employs sensors such as cameras and LiDAR to detect and model its surrounding environment, identifying both the target parking space and any nearby obstacles. Subsequently, a trajectory is planned based on the acquired perception data, and the vehicle’s speed and direction are continuously adjusted in real time—often via a PID controller or other conventional control algorithms to execute the parking maneuver. Although many of these algorithms have been successfully deployed in real-world vehicles, forming a mature industrial ecosystem, they remain highly dependent on accurate environmental information and precise vehicle dynamics models. As a result, perception errors can be amplified as they propagate through planning and control, increasing the risk of suboptimal maneuvers in complex or occluded environments.
Inspired by the Transformer model[vaswani2017attention], this paper proposes a pure vision-based end-to-end autonomous driving approach, effectively mitigating the aforementioned problems. Prior end-to-end parking pipelines often feed x- and y-coordinates alternately into a single decoder, which stretches the sequence and breaks the simultaneity of $(x_{t},y_{t})$ that holds in real vehicle motion, thereby accumulating errors. We therefore employ two decoders that produce $x_{t}$ and $y_{t}$ at the same time step and let them interact through Dual-Stream self-attention, yielding shorter sequences and stronger xxyy coupling. The specific contributions of this work are as follows:
@@
Figure 1: Overall architecture. Given the user-selected slot and multi-camera images, we form a Gaussian soft query and build BEV features via EfficientNet+LSS. Dual decoders predict $(x_{t},y_{t})$ with Dual-Stream Interactive Self-Attention, and a Soft Localization head refines each step using local BEV evidence before outputting the final trajectory (used by the controller).
@@ II RELATED WORK
II-A BEV representations for autonomous driving
Bird’s-Eye View (BEV) representations offer a clear, top-down perspective, showing vehicles, roads, and obstacles at their true planar positions. Consequently, BEV is widely adopted in autonomous driving applications. However, cameras cannot directly acquire distance information for each pixel. To address this, LSS [philion2020lift] compensates for the lack of depth information by predicting the depth distribution of each pixel. BEVDepth [li2023bevdepth], on the foundation of LSS, incorporates explicit depth supervision and a Depth Refinement Module, enhancing the reliability of depth estimation. BEVDet[huang2021bevdet] was improved on the basis of LSS and achieved higher accuracy. BEVDet4D [huang2022bevdet4d], built on BEVDet, introduces a time dimension to fuse spatiotemporal features. BEVFormer[li2022bevformer] is a Transformer-based BEV feature generation algorithm. It can simultaneously integrate the spatial and temporal information of multi-view images and reduce the computational load, generating high-quality features. DualBEV[li2024dualbev] introduces a dual feature fusion strategy to effectively predict BEV probabilities. BEVFusion[liu2023bevfusion] implements a multimodal fusion perception scheme that combines LiDAR and image data, effectively preserving both geometric and semantic information. FastBEV [li2024fast] proposed Fast-Ray, thereby significantly enhancing the inference speed. These BEV methods motivate our choice of LSS to obtain stable camera-to-BEV features. Unlike the binary slot query used in ParkingE2E [li2024parkinge2e], we adopt a Gaussian soft query that provides continuous spatial weights and smoother gradients within the target region (see Fig. 2).
@@ II-B End-to-End Autonomous Driving
In recent years, neural networks have substantially advanced end-to-end autonomous driving [bojarski2016end]. For clarity, we group prior work into city-driving and parking scenarios and review representative approaches in each.
Urban driving: Early end-to-end approaches directly map front-view images and commands to controls [codevilla2018end], while later methods inject trajectory supervision [wu2022trajectory] or fuse multimodal inputs with transformers [chitta2022transfuser, jaeger2023hidden, shamsoshoara2024swaptransformer, mu2024pix2planning, lyu2024sensor]. Despite strong results in city driving, many pipelines either predict controls directly or predict 2D waypoints/trajectories, and are not specifically tailored to the low-speed, high-precision geometric constraints of parking.
Parking: Parking-focused works include two-stage optimization [chai2018two], reinforcement learning [zhang2019reinforcement], and end-to-end pipelines that map images to controls or trajectories [shen2020parkpredict, shen2022parkpredict+, chai2022deep, yang2024e2e, li2024parkinge2e]. We follow the “images $\rightarrow$ trajectories” route but differ from prior work by (i) synchronously decoding $(x_{t},y_{t})$ with dual decoders, (ii) enabling explicit cross-stream interaction, and (iii) applying a soft-localization refinement that leverages local BEV context before emitting the final coordinates.
Most relevant baseline: ParkingE2E [li2024parkinge2e] generates a binary map around the selected slot and convolves it to form a query vector that attends to BEV features to predict trajectory points. While this explicit spatial prior highlights the target area, it offers no gradation within the region and creates sharp boundaries that may reduce attention continuity. We replace it with a Gaussian soft query and further add a soft-localization refinement head (see Fig. 2 and Fig. 4).
@@ III-A Problem Description and Optimization Goals
$\begin{gathered}\theta^{*}=\underset{\theta}{\operatorname*{arg\,min}}\,\mathbb{E}_{\left(I,\mathcal{X}^{\prime},\mathcal{Y}^{\prime},S,\mathcal{M}\right)\sim\mathcal{D}^{\prime}}\bigl[\mathcal{L}_{x}\bigl(\mathcal{X}^{\prime},\mathcal{N}^{x}_{\theta}(I,S)\bigr)\\ \quad+\mathcal{L}_{y}\bigl(\mathcal{Y}^{\prime},\mathcal{N}^{y}_{\theta}(I,S)\bigr)+\mathcal{L}_{m}\bigl(\mathcal{M},\mathcal{N}^{p}_{\theta}(I,S)\bigr)\bigr],\end{gathered}$ (6)
@@ III-B End-to-End Neural Network Model Based on Camera
Query: In ParkingE2E[li2024parkinge2e], the authors first generate a binary map with the same spatial dimensions as the BEV feature map and a depth of 1. In this map, the target slot and its surrounding fixed region are marked with a value of 1, while the area outside this range is marked with 0. After applying convolutional neural network, this map is used as a query vector to compute attention with the BEV feature map, serving as the final output of the encoder. This approach offers us significant inspiration and insight. The method of ParkingE2E clearly indicates to the model which areas need attention, However, a drawback is that it does not distinguish subregions within the target area, and the abrupt boundary transition might adversely affect continuity in attention. Therefore, in our work, we aim to give greater weight to the center while retaining some information from the surrounding areas, and this weight gradually weakens as the distance increases. We can select the desired parking slot by using the interactive tool ‘2D Nav Goal’ of Rviz and send the coordinates to the model. Subsequently, a zero tensor with the same width and height as the BEV feature map and a depth of 1 is created. In this channel, a probability map following a Gaussian kernel distribution is generated within the specified range centered on the coordinates of target slot. Subsequently, the tensor undergoes feature extraction via same convolutional neural network, transforming it into a high-dimensional vector. The formula is as follows:
@@
Key and Value: We employ EfficientNet [tan2019efficientnet] to extract image features $F_{img}\in\mathbb{R}^{C\times H_{img}\times W_{img}}$ from the RGB input, and then use the LSS[philion2020lift] method to generate the BEV feature map. First, we predict the depth distribution for each pixel and map it to the 3D space. Then, we fuse the depth information with the image features. Finally, we project the fused features onto the BEV voxel grid using the camera extrinsics and intrinsics to generate the camera features. For the obtained feature map, the range in the $x$-direction is denoted as $R_{x},R_{x}]$ m, where m indicates meters, and the range in the $y$-direction is denoted as $R_{y},R_{y}]$ m.
@@
We separate the X and Y coordinates of the model in the ParkingE2E [li2024parkinge2e], and then train and predict using two decoders with the same number of layers. At each time step, the model simultaneously outputs the X and Y coordinates, and they can only see the historical information of the same type of coordinates. Based on the findings of the comparative experiments depicted in Table I, the experimental outcome of jointly determining the most appropriate X-coordinate and Y-coordinate at one time based on the perceived surrounding environmental information surpasses the approach of splitting the X-coordinate and Y-coordinate into two time steps.
@@
To address the deficiency that the X and Y coordinates have no knowledge of each other’s historical information, we propose to initially concatenate the feature vectors of the X and Y coordinates at each time step, then conduct a self-attention mechanism calculation and obtain the fused feature vector. Subsequently, the fused feature vector is separated into x and y, which are respectively input into their respective decoders. The specific structure is presented as shown in the Figure 3. This attention module ensures that the high-dimensional feature vectors of the X and Y coordinate at the same time step are capable of integrating each other’s feature information. It guarantees that, in the subsequent mask self-attention layers in their respective decoders, The $X_{t}$ and $Y_{t}$ can indirectly take into account the X-coordinate and Y-coordinate information from 0 to t1t-1.
@@ III-C Lateral control and longitudinal control
In our work, we adopt the similar approach as ParkingE2E[li2024parkinge2e] for Lateral and Longitudinal Control of non-omnidirectional vehicle. During the control process, we mark the initial moment of parking as $t_{0}$. The predicted path $\mathcal{X}_{t_{0}},\mathcal{Y}_{t_{0}}=\mathcal{N}_{\theta}(I_{t_{0}},S)$ obtained at the initial moment is used. The relative pose from the initial moment $t_{0}$ to the current moment t can be obtained through the positioning system, denoted as $ego_{t_{0}\rightarrow t}$. The target steering angle $\mathcal{A}^{tar}$ is obtained using the RWF (Rear-wheel Feedback) method, and its expression is as follows:
@@
Figure 5: The inference results in indoor and outdoor scenarios. The orange line represents the inference trajectory of the method proposed in this paper, and the purple line represents that of ParkingE2E [li2024parkinge2e].
@@ IV-A The construction of the dataset
The dataset of this work originates from the publicly available real-vehicle dataset offered by ParkingE2E[li2024parkinge2e], which does not contain sudden obstacles and negative samples. This dataset acquires RGB images around the vehicle through on-board cameras and employs sensor data fusion algorithms to achieve precise vehicle positioning. Meanwhile, it provides the two-dimensional coordinates of the parking target slot in the ego coordinate system for each time step.
@@ IV-C Comparative experiment
To guarantee the fairness of the comparative experiments, the basic training parameters employed by all the models remain consistent. The size of the BEV Features is $200\times 200$, The actual spatial range is $x\in[-10\mathrm{m},10\mathrm{m}]$, $y\in[-10\mathrm{m},10\mathrm{m}]$ with a resolution of $0.1$ meters. In decoders, the maximum value of the trajectory serialization $N_{t}$ is $1200$. The trajectory decoder generates a sequence of predictions with a length of $30$. Our training equipment: One NVIDIA GeForce RTX 4090 GPU. The experimental outcomes of the ParkingE2E model[li2024parkinge2e] proposed in previous works, the transfuser model (based on the GRU decoder)[chitta2022transfuser], and our proposed model are presented in Table II. Under the same dataset and settings, our approach attains 0.2156/0.06296/1.239 on Hausdorff/L2/Fourier, compared with 0.386/0.1193/2.431 for ParkingE2E (Table II).
@@
ParkingE2E[li2024parkinge2e] 0.386 0.1193 2.431
TransFuser[chitta2022transfuser] 0.5293 0.3517 8.674
Figure 6: The performance of the two algorithms in the same scenario where an obstacle suddenly appears. The orange line represents the inference trajectory of the method proposed in this paper, and the purple line represents that of ParkingE2E[li2024parkinge2e]. Each row of images shows the trajectory inference results of four time steps from left to right.
@@ IV-D Visualization and Analysis of inference
We visualize the trajectory of model inference in Rviz, as shown in Figure 5. We also demonstrate that in the absence of special scenarios such as the sudden appearance of fast-moving objects in the training set, if other vehicles or pedestrians suddenly approach rapidly during the parking process, the model will immediately change its trajectory to prevent a collision. However, the model of ParkingE2E [li2024parkinge2e] mainly focuses its attention on the target area, thus being unable to modify the vehicle’s driving trajectory in a timely manner, as shown in Figure 6. Meanwhile, we noticed that during the parking process, if the vehicle is still at a distance from the parking space and there are no obstacles, our model will first plan a relatively straight trajectory to avoid turning too early and crossing the parking line. Only when the vehicle approaches the parking space will it start to plan a curve to achieve the turn. Moreover, with the same number of Predicted token, the trajectory inferred by ParkingE2E tends to have a larger range and sometimes deviates from the route, inferring a tortuous path. The specific effect is shown in Figure 7.
Figure 7: The performance of the two algorithms in the same indoor scene. The orange line represents the inference trajectory of the method proposed in this paper, and the purple line represents that of ParkingE2E[li2024parkinge2e].